HUBBLE SPACE TELESCOPE DAILY REPORT #1600 PERIOD COVERED: 0000Z (UTC) 03/27/96 - 0000Z (UTC) 03/28/96 Daily Status Report as of 088/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flat) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.2 Completed GHRS/2 5860 (Detection of ^11B/^10B and Galactic = Chemical Evolution) The High Resolution Spectrograph (Side Two) was used in the Echelle= Mode to make spectrographic observations of star HD102870 to study the = isotope-split Boron-11 and Boron-10 lines at 2090 A0. A wavelength = calibration observation was also taken. The observations were completed = as planned, and no problems were reported. 1.3 Completed WF/PC-2 6332 (The Circumnuclear Regions of Seyfert = Galaxies LRF Part: cycle 4 Medium) The WF/PC-2 and the linear ramp filter (LRF) were used to make = observations for the circumnuclear regions of Seyfert galaxy NGC-4388 in = both emission and continuum. From the data the PI expects to derive = sufficient evidence to confirm or refute a suggested model of Seyfert = Type 2 galaxies. The observations were executed as scheduled, and no = anomalies were noted. 1.4 Completed FOS/RD 6165 (FOS Cycle 5: Discriminator Test) This is a calibration proposal, which is being used to verify the = optimal discriminator settings for the Faint Object Spectrograph. The = observations were executed as scheduled, and no anomalies were noted. 1.5 Completed FOS/RD 6089 (Eclipse Mapping and Tomography of the UV = Emitting Region of DQ Her) The Faint Object Spectrograph (Red Side) was used to make rapid = mode ultraviolet observations of the eclipsing binary star 14Aur-C in = order to do eclipse mapping and tomographic studies. Observations were = made of both the ultraviolet continuum and emission lines. The = observations were executed as scheduled, and no anomalies were noted. 1.6 Completed FOS/RD 6167 (FOS Cycle 5: Dark Monitoring) The Faint Object Spectrograph (Red Side) was used to make several = dark calibration images. All exposures were internals. All spectra was = done in the image mode. The observations were completed as planned, and = no problems were reported. 1.7 Completed WF/PC-2 6344 (Low - Mass Binaries in the Hyades - = Completing the Survey) The WF/PC-2 was used to make observations of the low mass binary = star RHY077 (M-Class Dwarf Stars) which is located in the Hyades Cluster. = The FHST update that occurred just before this proposal failed, but no = effect on the observations was reported. No further anomalous activity = was noted. 1.6 Completed WF/PC-2 6102 (Chemical Evolution and History of Star = Formation in the Large Magallanic Cloud) The WF/PC-2 was used to make observations of the "Bar" region in = the Large Magellanic cloud to study the chemical evolution and the = history of star formation in that region. The goal is to observe = individual field stars near the main sequence turnoff in order to study = the large scale processes that control star formation and nucleosynthesis = in galaxies. The observations were executed as planned, and no anomalies = were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 7 As documented in HSTAR 5551, the acquisition at 088/073646Z failed = to guidance on FGS #3 only. Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 19 Successful : 18 The update at 087/1523Z failed. 2.3 Operations Notes: Due to maintenance work being performed in the MOR/SMOR area, HST = operations transitioned to the SAMS backup facility yesterday from = approximately 8:00 a.m. until 4:00 p.m. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH